OpenStudy (anonymous):

Simplify the radical expression. Sqrt4x^4

TheSmartOne (thesmartone):

\[\sqrt{4x ^{4}}\]

TheSmartOne (thesmartone):

Is that your question?

TheSmartOne (thesmartone):

@adoredelano

TheSmartOne (thesmartone):

So you can make it \[\sqrt{4x ^{4}}=\sqrt{4}\times \sqrt{x^4}=2\times x^2=2x^2\]